OpenStudy (kkern00):

use interval notation to express the solution set and graph the solution set on a number line. 6 - (x+5) > 1 - 3x ??

satellite73 (satellite73):

from logs to inequalities? what is this, on line kitchen sink math course?

satellite73 (satellite73):

distribute the minus sign on the left hand side first what do you get?

OpenStudy (kkern00):

@satellite73 math assessment for college lol

OpenStudy (kkern00):

@satellite73 6 - x - 5 > 1 - 3x

satellite73 (satellite73):

yes right

satellite73 (satellite73):

now combine like terms, i.e. compute on the left \(6-5\) what do you have then ?

OpenStudy (kkern00):

just 1

OpenStudy (kkern00):

i did get to x - 1 > -3x +1 but i was lost from there on

satellite73 (satellite73):

actually it would not be that, it would be \[1-x>-3x+1\]

satellite73 (satellite73):

add \(3x\) to both sides

OpenStudy (kkern00):

1 + 2x > 1 ?

satellite73 (satellite73):

yes, now subtract 1 from both sides

OpenStudy (kkern00):

2x > 0 ?

satellite73 (satellite73):

right now you can either say "done" or you can divide both sides by 2, still get zero on the right

OpenStudy (kkern00):

these two graphs look insanely similar I'm not sure of the difference?

satellite73 (satellite73):

they sure do

satellite73 (satellite73):

but one difference, the round vs the square bracket

OpenStudy (kkern00):

but i am thinking A?

satellite73 (satellite73):

round for \(x>0\) square for \(x\geq 0\)

OpenStudy (kkern00):

Right i saw that

satellite73 (satellite73):

so if you have \[x\geq 0\] go for A if \[x>0\] go for B

OpenStudy (kkern00):

it was technically \[x \ge\]

OpenStudy (kkern00):

i just didn't have that option typing out the question

satellite73 (satellite73):

yeah that is why i did not pick , figured it could be either

OpenStudy (kkern00):

thank you!

satellite73 (satellite73):

yw